# Robust Adaptive Safe Robotic Grasping with Tactile Sensing

### v1

@@ VI-A Experimental Setup
The proposed framework is developed and implemented in C++ via an ASUS FX507ZC4 laptop with i7-12700H processor, 16GB of RAM, and NVIDIA GeForce RTX 3050 GPU. Our framework operates as a high-level controller at 125Hz while a low-level tracking controller runs at 1k Hz via an Intel NUC with an i7-10710U processor and 4GB of RAM. The implementation of quadratic programming in safety filters is carried out using QuadProg++ library [36] which is based on [37]. The objects used in this experiment include a small wooden cube (10g), a plastic cylinder (30g), an aluminium bearing housing (120g), and a thin, fragile piece of lab glassware (15g) as shown in the first column of Fig. 6. The friction coefficients $\hat{\mu}$ for each object are iteratively tuned to correspond to the minimum contact normal force required to grasp a given object without slippage.
As shown in the middle column of Fig. 6, we intentionally generate the desired force (the dashed red line) that violates safety constraints in order to evaluate the performance of each safety filter. For the safety constraints, we define a force regulation constraint for the contact normal force, $h_{1}$. The minimum normal reaction forces in this experiment are chosen as $-3$N, $-3$N, $-3.5$N, and $-2.5$N in the opposite direction to the applied force for the cube, the cylinder, the housing, and the glass object, respectively. For force closure constraints, we consider only a vertical cone plane with respect to the gravity direction, denoted by $h_{2}$ and $h_{3}$ across all experiments since the finger does not move in the horizontal direction at the contact point. Lastly, we use a linear function for an extended class $\mathcal{K}_{\infty}$ function, $\alpha(\cdot)$, (e.g., $\alpha h(\bm{x})$). The $\alpha$ values are the same for each safety filter, but vary for each object. For example, we conduct the experiment with a cube with $\alpha=(80,65,65)$ corresponding to each constraint, $h_{1},h_{2},h_{3}$, respectively. The rest of the experiment parameters are presented in Table I.
@@ VI-B Results and Discussion
Figure 6 presents the overall experimental results of our proposed framework. As shown in the middle column of Fig. 6 (a), the standard CBF (blue) failed to maintain force regulation and closure constraints due to the model uncertainty, resulting in excessive force applied to the object at first, followed by slippage due to insufficient grasping force at the end. We also observed that $h_{3}$ became negative as shown in the last column of Fig. 6 (a), indicating a safety violation. In contrast, the RCBF (green) showed significantly more robustness to uncertainty due to the maximum disturbance bound in the safety constraints. Likewise, we observe that the RaCBF is slightly robust to model uncertainty, but exhibits less conservative behavior than the RCBF since it adapts to the parametric uncertainty in the safety conditions. While RaCBF still has conservatism, the DOBCBF shows the least conservative performance since the estimated disturbances are directly attenuated in the safety conditions. Table I provides the minimum values of each safety condition for each safety filter, indicating that the CBF violates safety while other methods ensure safety.
In the cases of the plastic cylinder, the housing, and the fragile glassware, we observe that the standard CBF violates safety, while the other methods maintain safety as shown in Fig. 6 (b)-(d) and Table I. As indicated in Table I, the conservative behaviors of each safety filter increasingly improve in the following order (with the exception of the glassware): RCBF, RaCBF, and DOBCBF. Since the feasible safe force region for the fragile object is much smaller than other objects, it is shown that the performance of each safety filter is similar in terms of conservatism. Compared to [29] using RCBF only, the proposed method leverages RaCBF and DOBCBF, thereby reducing conservatism by $83.5\%$ on average.
Despite these positive results, numerous challenges were encountered in the hardware trials. First, there still exists uncertainty over $\mu$ even after iteratively tuning $\hat{\mu}$ to satisfy the force closure condition. To address this concern, we could employ friction coefficient estimation methods [38] from prior work for grasping to further reduce uncertainty in the force closure condition. Second, there is uncertainty in the contact frame on a grasped object. In the experiments, we assume that the contact frame is approximated as the finger’s frame, which leads to model uncertainty in the system, deteriorating our safety guarantees. However, these uncertainties can still be mitigated by using robust approaches such as RaCBFs, RCBFs or DOBCBFs, as demonstrated in our experiments. Another approach is to integrate in-hand pose estimation [39] to reduce the uncertainty for the contact frame, which can enhance the performance of the proposed framework.
@@ VII Conclusion
This paper presented a comprehensive framework for safe grasping with formal safety guarantees by using CBFs. We implemented three main components of the framework: safety filters, finger force controller, and the estimation of contact point and force/torque. We first designed contact force and force closure constraints and then enforced them into various safety filters to achieve safe grasping. We implemented and developed a finger force controller to track safe control input and a technique to estimate contact point and force/torque from multiple tactile sensors at each finger. We simulated the framework in a scenario where two fingers grasp a cube, showing that the framework can achieve safe grasping with each safety filter. Finally, the proposed framework was experimentally validated with multiple objects including a fragile object on Shadow Robot setup, resulting in successful safe grasping without safety violations across all experiments. In the simulation and experiments, we also evaluated the performance of each safety filter in terms of safety violation and conservatism, and found that using DOBCBF is significantly advantageous for achieving safe grasping, since it provides the best performance to ensure safety guarantees with minimum conservatism.

### v2

@@ VI-A Experimental Setup
The proposed framework is developed and implemented in C++ via an ASUS FX507ZC4 laptop with i7-12700H processor, 16GB of RAM, and NVIDIA GeForce RTX 3050 GPU. The framework operates as a high-level controller at 125Hz while a low-level tracking controller runs at 1k Hz via an Intel NUC with an i7-10710U processor and 4GB of RAM. The implementation of quadratic programming in safety filters is carried out using QuadProg++ library [Liu2021] which is based on [Goldfarb1983]. The objects used in this experiment include a small wooden cube (10g), a plastic cylinder (30g), an aluminium bearing housing (120g), and a thin, fragile piece of lab glassware (15g) as shown in the first column of Fig. 6. The friction coefficients $\hat{\mu}$ for each object are iteratively tuned to correspond to the minimum contact normal force required to grasp a given object without slippage.
As shown in the middle column of Fig. 6, we intentionally generate the desired force (the dashed red line) that violates safety constraints in order to evaluate the performance of each safety filter. For the safety constraints, we define a force regulation constraint for the contact normal force, $h_{1}$. The minimum normal reaction forces in this experiment are chosen as 3-3N, 3-3N, 3.5-3.5N, and 2.5-2.5N in the opposite direction to the applied force for the cube, the cylinder, the housing, and the glass object, respectively. For force closure constraints, we consider only a vertical cone plane with respect to the gravity direction, denoted by $h_{2}$ and $h_{3}$ across all experiments since the finger does not move in the horizontal direction at the contact point. Lastly, we use a linear function for an extended class $\mathcal{K}_{\infty}$ function, $\alpha(\cdot)$, (e.g., $\alpha h(\bm{x})$). The $\alpha$ values are the same for each safety filter, but vary for each object. For example, we conduct the experiment with a cube with $\alpha=(80,65,65)$ corresponding to each constraint, $h_{1},h_{2},h_{3}$, respectively. The rest of the experiment parameters are presented in Table I.
@@ VI-B Results and Discussion
Figure 6 presents the overall experimental results of the proposed framework. As shown in the middle column of Fig. 6 (a), the standard CBF (blue) failed to maintain force regulation and closure constraints due to the model uncertainty, resulting in excessive force applied to the object at first, followed by slippage due to insufficient grasping force at the end. We also observed that $h_{3}$ became negative as shown in the last column of Fig. 6 (a), indicating a safety violation. In contrast, the RCBF (green) showed significantly more robustness to uncertainty due to the maximum disturbance bound in the safety constraints. Likewise, we observe that the RaCBF is slightly robust to model uncertainty, but exhibits less conservative behavior than the RCBF since it adapts to the parametric uncertainty in the safety conditions. While RaCBF still has conservatism, the DOBCBF shows the least conservative performance since the estimated disturbances are directly attenuated in the safety conditions. Table I provides the minimum values of each safety condition for each safety filter, indicating that the CBF violates safety while other methods ensure safety.
In the cases of the plastic cylinder, the housing, and the fragile glassware, we observe that the standard CBF violates safety, while the other methods maintain safety as shown in Fig. 6 (b)-(d) and Table I. As indicated in Table I, the conservative behaviors of each safety filter increasingly improve in the following order (with the exception of the glassware): RCBF, RaCBF, and DOBCBF. Since the feasible safe force region for the fragile object is much smaller than other objects, it is shown that the performance of each safety filter is similar in terms of conservatism. Compared to [Shaw2021] using RCBF only, the proposed method leverages RaCBF and DOBCBF, thereby reducing conservatism by $83.5\%$ on average.
Despite these positive results, numerous challenges were encountered in the hardware trials. First, there still exists uncertainty over $\mu$ even after iteratively tuning $\hat{\mu}$ to satisfy the force closure condition. To address this concern, we could employ friction coefficient estimation methods [WU2024114249] from prior work for grasping to further reduce uncertainty in the force closure condition. Second, there is uncertainty in the contact frame on a grasped object. In the experiments, we assume that the contact frame is approximated as the finger’s frame, which leads to model uncertainty in the system, deteriorating safety guarantees. However, these uncertainties can still be mitigated by using robust approaches such as RaCBFs, RCBFs or DOBCBFs, as demonstrated in the experiments. Another approach is to integrate in-hand pose estimation [InHandPoseEstimation] to reduce the uncertainty for the contact frame, which can enhance the performance of the proposed framework.
@@ VII Conclusion
This paper presented a comprehensive framework for safe grasping with formal safety guarantees by using CBFs. We implemented three main components of the framework: safety filters, finger force controller, and the estimation of contact point and force/torque. We first designed contact force and force closure constraints and then enforced them into various safety filters to achieve safe grasping. We implemented and developed a finger force controller to track safe control input and a technique to estimate a contact point and force/torque from multiple tactile sensors at each finger. We simulated the framework in a scenario where two fingers grasp a cube, showing that the framework can achieve safe grasping with each safety filter. Finally, the proposed framework was experimentally validated with multiple objects including a fragile object on Shadow Robot setup, resulting in successful safe grasping without safety violations across all experiments. In the simulation and experiments, we also evaluated the performance of each safety filter in terms of safety violation and conservatism, and found that using DOBCBF is significantly advantageous for achieving safe grasping, since it provides the best performance to ensure safety guarantees with minimum conservatism.